OpenStudy (anonymous):

A train traveled 300 miles. How long did the trip take if the train was traveling at a rate of 50 mph?

hartnn (hartnn):

time=distance /speed does this help ?

OpenStudy (zzr0ck3r):

d = r*t d = 300, r = 50, what is t?

OpenStudy (anonymous):

so t should = 6 ..right ?

OpenStudy (anonymous):

The basic equation is :- Speed(S) = Distance(D) / Time(T) S = D/T This can be algebraically re-arranged:- D = S x T Distance = Speed x Time & T = D/S Time = Distance / Speed

hartnn (hartnn):

right t=6 hours

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

what if the rate is 70 mph? I tried the same process with that but it came out weird

OpenStudy (zzr0ck3r):

t = d/r so d = 300 and r = 70 what did you get?

OpenStudy (zzr0ck3r):

its not going to be a nice number like 6

OpenStudy (anonymous):

yeah I noticed lol should I just round the number I received?

OpenStudy (zzr0ck3r):

what is it?

OpenStudy (anonymous):

I divided 300 by 70 and got this

OpenStudy (zzr0ck3r):

note if is hours your looking for and you get something like 6.7 you would want to find out what .7 of 1 hour is (7/10)*60 = 7*6 = 42 so 6.7 would be 6 hours and 42 minutes

OpenStudy (zzr0ck3r):

right, so you have 4.28 hours. find out what .28 of an hour is

OpenStudy (zzr0ck3r):

.28*60 = (28/100)*60 = ?

OpenStudy (zzr0ck3r):

I dont know if your teacher wanted you to do this, but its a good excercise. you could just call it 4.28 hours. but giving a clean answer would be better imo.

OpenStudy (anonymous):

well I do school online.. and so they don't show us much. I would think either answer would be fine

OpenStudy (anonymous):

I'm going to try the clean answer

OpenStudy (zzr0ck3r):

i would think so.

OpenStudy (zzr0ck3r):

ok tell me what you get

OpenStudy (anonymous):

it keeps giving me 16.8

OpenStudy (zzr0ck3r):

so 16.8 minutes so we are looking at about 4 hours and 16 minutes.

OpenStudy (zzr0ck3r):

you could find out how much of a min is .8 if you wanted and give it down to the second, but I think thats going a little to far:)

OpenStudy (anonymous):

alright, that's what I thought it would be

OpenStudy (anonymous):

lol yeah I don't my school cares about the seconds ..so 4 hours and 16 minutes it is ^.^

OpenStudy (anonymous):

on one they gave me x mph for a rate.. should x count as 1? How would I do that ?

OpenStudy (zzr0ck3r):

can you show the exact question?

OpenStudy (anonymous):

A train traveled 300 miles. How long did the trip take if the train was traveling at a rate of x mph? its the same question with a different rate

OpenStudy (zzr0ck3r):

when the rate was 50 we said t = 300/50 when the rate was 70 we said t = 300/70 when the rate is x we say t = 300/?

OpenStudy (anonymous):

300/x ? I don't understand how to work that problem

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

its just a variable x can be anything.

OpenStudy (anonymous):

so should the answer just be x ?

OpenStudy (zzr0ck3r):

no you said it already

OpenStudy (zzr0ck3r):

if x = 50 its 300/50 if x = 234567864 then its 300/234567864 if x = x then its 300/x we do not have a value for x. so it acts as a place holder. Just like if I said y = 54x+2 if x = 2 its 45(2)+2 if x = 54 its 45(54)+2 if x=x(or we are given no value for x) its 45x+2

OpenStudy (anonymous):

well for the answer I type in how long the trip would take.. for this should i just put x ? would that be a correct answer?

OpenStudy (zzr0ck3r):

no t = 300/x so time = 300/x

OpenStudy (anonymous):

because the next one is (x+20)mph ..so i would assume that would be 20x

OpenStudy (zzr0ck3r):

i must go to sleep time will always be t = distance/rate what ever they give you for the rate put that where the word rate is so if the rate was 45x+7654y+43243+dog+cat then you would have t = 300/(45x+7654y+43243+dog+cat )

OpenStudy (anonymous):

okies :) I understand now ^.^ thank you soo much for helping me :)